If john gets 97 on his next math test, his average will be 90. if he gets 73, his average will be 87. the number of tests that john has already taken is
A 6 b 7 c 8 d 9 e none of these

Answer by checkley77(12569)

x+97/(y+1)=90 cross multiply.
x+97=90y+90
x+73/(y+1)=87 cross multiply.
x+73=87y+87 subtract the other equation.
-(x+97=90y+90)
------------------
-24=-3y-3
-3y=-24+3
-3y=-21
y=-21/-3
y=7 total tests previously taken befre the 90 or 73 grade test.